You are a nurse practitioner working in an ambulatory clinic in Illinois with a case load of 350 patients. You notice that a large proportion (179 patients) have COPD, and you wonder if it is related to smoking. You create the following 2 x 2 table to explore the relationship between smoking and COPD:

With COPD

Without COPD

total

Smokers

82

97

179

Non-Smokers

18

153

171

100

250

350

1. What is the rate of COPD among your patients who smoke? 82

2. What is the relative risk of COPD among the smokers and nonsmokers? 82/179= 0.46

18/171=0.11      0.46/0.11= 4.18

3. What is the risk for COPD that can be attributed to smoking?

4. What are the odds of getting COPD if a patient is a smoker? 82/18= 4.56

5. You learn that the rate of COPD among nonsmokers in the state of Illinois is 0.26. Why does that statistic cause you concern? What is your next step?

6. What other statistics would help you determine how big the problem is among your patients
Answer #1

1. rate of COPD among your patients who smoke : 82/179 = (approx) 0.458

2. proportion of COPD among smokers= 82/179

proportion of COPD among non-smokers= 18/171

Relative risk= (82/179)/(18/171) =(approx) 4.35

3. risk for COPD that can be attributed to smoking = 82/179 = (approx) 0.458

4.odds of getting COPD if a patient is a smoker= 82/97= (approx) 0.845

5.If the rate of COPD among nonsmokers in the state of Illinois is 0.26, then out of 100 people 26 people suffer from COPD which is quite high. Then we can think that other attributes are also present which may cause COPD.

6. We can check for the independence of the attributes smokers and COPD using chi-square test.